OpenStudy (anonymous):

Triangles ABC and DEF are congruent. If mA = mD = 48º, and mB = 67º, which is mC + mE? 134º 132º 130º 136º

myininaya (myininaya):

i think 132

OpenStudy (anonymous):

mA=mD=48 as triangle is congruent hence mC=mF=67 also mB=mE you can find mB=180-(48+67)=65 than=mC+mE=mC+mC =65+65=130 (because mC=mE)

OpenStudy (anonymous):

thank you!

myininaya (myininaya):

i thought angle B is 67 not 65

OpenStudy (anonymous):

so is it still 130?

myininaya (myininaya):

this is what i thought A=D=48 B=67 C=F B=E C=180-48-67=65 E=67 C+E=65+67=132